Nouabalé Ndoki (10.01.2014 to 12.31.2014)

You are here

PDF versionPDF version
Year: 
2014
Quarter: 
October - December
Communications activities: 
None
Unusual Events at the TEAM Site: 
None for now, data still being processed
New Species: 
None for now, data still being processed
Protocol Activities: 
As scheduled, the vegetation sampling and camera trap deployment occurred in Array 2 in December through to January 17, 2014 when cameras traps were collected. VG data were collected in the 3 plots in Array 2. Data are being processed and once ready, will be upload into the TEAM portal. Only 29 CT were deployed and there are just collected from the field. Then all process will be follow before uploading data into the TEAM portal CL data are being collected monthly and uploaded on the TEAM portal
Protocol Problems: 
During this camera trap deployment in Array 2, one camera encountered a technical problem. Indeed, although all the cameras were tested before leaving to the field, one of them did not start up once in the field. Given that we had no camera in reserve it was not possible to change it. Until now are still waiting for the new material which still in NY since September. It seems like there will be the same problem in Array 1 because the mission will be carried out in two weeks time
Schedule Problems: 
None
Logistical Problems: 
None
Suggestions: 
None